# Transition Path Sampling with Improved Off-Policy Training of Diffusion Path Samplers

### v4

@@ Abstract
Understanding transition pathways between meta-stable states in molecular systems is crucial to advance material design and drug discovery. However, unbiased molecular dynamics simulations are computationally infeasible due to the high energy barriers separating these states. Although recent machine learning techniques offer potential solutions, they are often limited to simple systems or rely on collective variables (CVs) derived from costly domain expertise. In this paper, we introduce a novel approach that trains diffusion path samplers (DPS) for transition path sampling (TPS) without the need for CVs. We recast the problem as an amortized sampling of the target path measure, minimizing the log-variance divergence between the path measure induced by our DPS and the target path measure. To ensure scalability for high-dimensional tasks, we introduce (1) a new off-policy training objective based on learning control variates with replay buffers and (2) a scale-based equivariant parameterization of the bias forces. We evaluate our approach, coined TPS-DPS, on a synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin. Results show that our approach produces more realistic and diverse transition pathways compared to existing baselines. We also provide links to our project page and code.
@@ 1 Introduction
In material design and drug discovery, it is crucial to understand the mechanisms and kinetics of transitions between meta-stable states of molecular systems, such as protein folding and chemical reactions (Mulholland, 2005; Piana et al., 2012; Ahn et al., 2019; Spotte-Smith et al., 2022). Their comprehensive study requires sampling transition paths (Elber, 2016; Lee et al., 2017), which provides insight into mechanisms and energy landscapes. However, naïvely sampling transition paths by molecular dynamics (MD) simulations is often computationally intractable due to high energy barriers, which cause an exponential decay in probability to make a transition (Pechukas, 1981).
@@
To address this problem, researchers have developed enhanced sampling approaches such as steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999), umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), and adaptive biasing force (ABF; Comer et al., 2015) methods. These methods rely on bias forces to facilitate transitions across high energy barriers. They are mainly designed based on collective variables (CVs), which are functions of atomic coordinates that capture the slow modes of the transition. Although effective for some systems, the reliance on expensive domain knowledge limits the applicability of the methods to systems where CVs are less understood.
@@
Recently, machine learning has emerged as a promising paradigm for CV-free transition path sampling (Das et al., 2021; Lelièvre et al., 2023; Holdijk et al., 2024). The key idea is to parameterize the bias force using a neural network and train the corresponding biased MD simulation to sample realistic transition paths. In particular, Lelièvre et al. (2023) considered reinforcement learning to sample paths escaping meta-stable states. Das et al. (2021) and Holdijk et al. (2024) proposed to minimize the Kullback-Leibler (KL) divergence between the path measures induced by the neural network and the target path measure. However, minimizing the KL divergence leads to mode collapse, which is hazardous for sampling transition paths due to a lack of exploration. Furthermore, Das et al. (2021) and Lelièvre et al. (2023) limited their evaluation to low-dimensional synthetic systems. Designing machine learning algorithms for CV-free transition path sampling of large molecules remains an open problem.
Contribution. In this work, we propose to solve the transition path sampling problem with the diffusion path sampler (TPS-DPS). TPS-DPS trains the bias force to minimize a recently proposed log-variance divergence (Nüsken & Richter, 2021) between the path measure induced by the bias MD and the target path measure. To ensure scalability for high-dimensional tasks, e.g., Chignolin folding, we propose a new off-policy training scheme and a scale-based parameterization of the bias forces.
To be specific, we propose a new training objective to minimize the log-variance divergence, defined between a pair of path measures induced by the biased MD and the unbiased MD conditioned on the target meta-stable state. To minimize the variance of gradient estimators and improve sample efficiency, we propose to use (1) learnable control variates and (2) replay buffers with simulated annealing. Our choice of minimizing the log-variance is motivated by recent works (Richter et al., 2020; Nüsken & Richter, 2021; Richter & Berner, 2024) that demonstrate many desirable properties of the log-variance divergence such as a robust gradient estimator for training stability and avoidance of differentiation through the SDE solver. This allows the use of previously collected samples in a replay buffer, promoting sample efficiency in training and improving the diversity of the generated transition paths by preventing the mode collapse observed in KL divergence.
@@
We also introduce a new $SE(3)$ equivariant scale-based parameterization for the bias force that guides the biased MD to sample meaningful paths more frequently, based on inductive bias. Our key idea is to parameterize the bias force to yield a positive inner product with a direction to decrease the distance between the molecular state and the target meta-stable state, which guarantees the bias force to (locally) decrease the distance. We also use the Kabsch algorithm (Kabsch, 1976) to align the molecule and the bias force to the target meta-stable state, guaranteeing $SE(3)$ equivariance for better generalization across molecular states that are identical up to roto-translations.
@@
We extensively evaluate our method on the synthetic double-well potential and three peptides: Alanine Dipeptide, Polyproline Helix, and Chignolin. We compare TPS-DPS with prior ML approach (PIPS; Holdijk et al., 2024), as well as classical non-ML methods, e.g., steered MD (SMD; Izrailev et al., 1999). Our experiments demonstrate that TPS-DPS consistently generates realistic and diverse transition paths, similar to the ground truth ensemble. In addition, ablation studies of algorithmic components verify the effectiveness of our method.
@@ 2 Related work
Transition path sampling without ML. Metadynamics (Branduardi et al., 2012), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), adaptive biasing force (ABF; Comer et al., 2015), and steered molecular dynamics (SMD; Schlitter et al., 1994; Izrailev et al., 1999) were introduced to explore molecular conformations that are difficult to access by unbiased molecular dynamics (MD) within limited simulation times (Hénin et al., 2022). However, they mostly rely on collective variables (CVs) for high-dimensional problems and are inapplicable to systems with unknown CVs. To sample transition paths without CVs, Dellago et al. (1998) proposed shooting methods that use the Markov chain Monte Carlo (MCMC) procedure on path space.
@@
Data driven ML approaches. Recently, generative models have been trained to sample a new transition path given a dataset of transition paths. Petersen et al. (2023); Triplett & Lu (2023) and Lelièvre et al. (2023) applied diffusion probabilistic models (Ho et al., 2020) and variational auto-encoders (Kingma & Welling, 2013) for transition path sampling, respectively. However, these methods are limited to small systems. Concurrent to our work, Jing et al. (2024) generated molecular trajectories by task-wise conditions given an initial state and viewed transition path sampling as an interpolation task between two molecular states.
We also note some related works that are not directly comparable to our method. Klein et al. (2024) proposed to accelerate MD simulation through coarse-grained path generation that may include multiple meta-stable states. However, this work is incapable of targeting a specific meta-stable state and generating fine-grained transition paths. Duan et al. (2023); Kim et al. (2024) use neural networks to generate transition states of a given chemical reaction, but cannot generate transition paths.
@@
Data free ML approaches. Without a previously collected dataset, Das et al. (2021); Lelièvre et al. (2023); Hua et al. (2024); Holdijk et al. (2024) trained generators to sample transition paths using biased MD. Lelièvre et al. (2023) used reinforcement learning to train the generator but focused on escaping an initial meta-stable state rather than reaching the target meta-stable state. Next, Holdijk et al. (2024) considered transition path sampling as minimizing the reverse KL divergence between path distribution from biased MD and transition path distribution. In this work, we mainly compare our method with Holdijk et al. (2024) and provide quantitative analysis for real-world molecule systems. We also note that, concurrent to our work, Du et al. (2024) also considered variational formulation of the TPS problem using Doob’s $h$-transform.
@@ 3 Transition path sampling with diffusion path samplers
Figure 1: Problem setup. Potential energy landscape of Alanine Dipeptide, where the sampled path from state $\mathcal{A}$ to state $\mathcal{B}$ is highlighted in yellow dotted lines. We visualize snapshots in the transition path highlighted in white circles and the transition state highlighted in the white star.
In this section, we introduce our method, coined transition path sampling with diffusion path sampling (TPS-DPS). Our main idea is to formulate the transition path sampling (TPS) problem as a minimization of log-variance divergence (Nüsken & Richter, 2021) between two path measures: the path measure of transition paths and that of biased MD simulations parameterized by a neural network. Our main methodological contribution is twofold: (1) a new training algorithm that minimizes the log-variance divergence with a learned control variate and replay buffer, and (2) a $SE(3)$ equivariant scale-based parameterization of the biased MD that provides inductive bias for dense training signals in high-dimensional problems.
@@ 3.1 Problem setup
Our goal is to sample transition paths from one meta-stable state to another meta-stable state given a molecule system. We provide an example of the problem for Alanine Dipeptide in Figure 1. We view this as a task to sample paths from an unbiased molecular dynamics (MD), described by a stochastic differential equation (SDE), conditioned on its starting and ending points. To solve this task, we train the bias force parameterized by a neural network to amortize the sampling procedure.
Molecular dynamics. We consider a MD simulation on time interval $[0,T]$ that describes motion of a molecular state $\bm{X}_{t}=(\bm{R}_{t},\bm{V}_{t})\in\mathbb{R}^{6N}$ at time $t$ where $N$ is the number of atoms, $\bm{R}_{t}\in\mathbb{R}^{3N}$ is the atom-wise positions and $\bm{V}_{t}\in\mathbb{R}^{3N}$ is the atom-wise velocities. In particular, we adopt Langevin dynamics (Bussi & Parrinello, 2007) defined as the following SDE:
@@
where $U$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\bm{W}_{t}$ denote the potential energy function, the atom-wise masses, the friction term, the Boltzmann constant, the absolute temperature, and the Brownian motion, respectively, and $\bm{\zeta}\in\mathbb{R}^{3N}$ is a vector of positive infinitesimal values. We let $\mathbb{P}_{0}$ denote the path measure induced by Equation 1 in the space $\mathcal{C}([0,T];\mathbb{R}^{6N})$ of continuous functions $\bm{X}:[0,T]\rightarrow\mathbb{R}^{6N}$.
@@
To be specific, we aim to sample from the target path measure $\mathbb{Q}$ which reweights the path measure $\mathbb{P}_{0}$ via the Radon–Nikodym derivative, defined as follows:
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\frac{1_{\mathcal{B}}(\bm{X})}{\mathbb{E}_{\mathbb{P}_{0}}\left[1_{\mathcal{B}}(\bm{X})\right]},\quad 1_{\mathcal{B}}(\bm{X})=\begin{cases}1&\text{if}~{}\lVert\bm{R}_{T}-\rho_{T}\cdot\bm{R}_{\mathcal{B}}\rVert\leq\delta,\\ 0&\text{otherwise},\end{cases}$ (2)
where $\cdot$ denotes group action associated with the $SE(3)$ space and $\rho_{T}\cdot\bm{R}_{\mathcal{B}}$ is the aligned position by the optimal roto-translation $\rho_{T}\in SE(3)$ to minimize its Euclidean distance to $\bm{R}_{T}$, i.e., $\rho_{T}=\operatorname*{arg\,min}_{\rho\in SE(3)}\lVert\bm{R}_{T}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$. Such a transformation can be obtained from the Kabsch algorithm in $O(N)$ complexity (Kabsch, 1976).
Note that one may consider naïve rejection sampling to sample transition paths, based on running unbiased MD to sample a path $\bm{X}$ from the path measure $\mathbb{P}_{0}$ and accepting if the path $\bm{X}$ arrives at the neighborhood of the position $\bm{R}_{\mathcal{B}}$ with the radius $\delta$. However, this method does not scale to high-dimensional or low-temperature problems, since the sampled path by unbiased MD rarely reaches the target states due to the high energy barriers, i.e., the rejection ratio is too high.
@@ 3.2 Log-variance minimization
In this section, we propose our algorithm to amortize transition path sampling. Our key idea is to train a neural network to induce a path measure that matches the target path measure $\mathbb{Q}$, using the log-variance divergence (Nüsken & Richter, 2021) between the path measures. We propose a new training scheme to minimize the log-variance divergence based on learning the control variate of its gradient and a replay buffer to improve sample efficiency and diversity.
@@
The first three terms in Equation 6 correspond to the deviation of the biased MD from the unbiased MD integrated over the path sampled from $\mathbb{P}_{\tilde{\bm{v}}}$. The last term reweights the unbiased MD to the target path measure $\mathbb{Q}$. As a result, minimizing Equation 5 could be thought as minimizing the variation between $\mathbb{P}_{\bm{v}_{\theta}}$ and $\mathbb{Q}$. We provide the full derivation in Section A.1. Compared to KL divergence, the log-variance divergence provides a robust gradient estimator and avoids differentiating through the SDE solver. (Richter et al., 2020; Nüsken & Richter, 2021).
Minimizing with learnable control variate. To minimize the log-variance divergence, we consider the following loss function that replaces the estimation of $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ by learning a parameter $w$:
@@
where $w$ is a control variate that controls the variance of the gradient estimator of $\nabla_{\theta}\mathcal{L}(\theta,w)$ without changing the gradient. Note that we set $\tilde{\bm{v}}=\bm{v}_{\theta}$ in Equation 6, which implies that the gradient of Equation 7 coincides with the KL divergence (Richter et al., 2020; Nüsken & Richter, 2021). When optimized, the control variate $w$ estimates the expectation $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ since $\operatornamewithlimits{argmin}_{w}\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$. Thus, jointly optimizing $(\theta,w)$ with the gradient step can be interpreted as jointly minimizing log-variance divergence and estimating $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ using $w$.
Off-policy training with replay buffer. To leverage the degree of freedom in reference path measure for the log-variance divergence, we allow discrepancy between reference path measure and current path measure, called off-policy training, which is widely used in discrete-time reinforcement learning (Mnih et al., 2013; Bengio et al., 2021). For the sample efficiency, we reuse the samples with a replay buffer $\mathcal{D}$ which stores path samples from the path measure $\mathbb{P}_{\bm{v}_{\bar{\theta}}}$ associated with previous policies $\bm{v}_{\bar{\theta}}$. Our modified loss function $\mathcal{L}^{\mathcal{D}}$ with replay buffer $\mathcal{D}$ is defined as follows:
@@
Using the replay buffer also prevents mode collapse, using diverse paths from different path measures. Similar to other off-policy training algorithms (Malkin et al., 2022; Kim et al., 2023), we use simulated annealing to collect diverse paths that cross high energy barriers.
@@
We describe our training algorithm in Section A.3. Overall, our off-policy training algorithm iterates through four steps: (1) sampling paths from the biased MD simulation with current policy $\bm{v}_{\theta}$ at high temperature, (2) storing sampled paths in the replay buffer $\hat{\mathcal{D}}$, (3) sampling a batch of the paths from the replay buffer, and (4) training current policy $\bm{v}_{\theta}$ by minimizing the loss in Equation 10. After minimization, biased MD simulation can directly sample transition paths by amortizing inference in the target path measure.
3.3 Parameterization for high dimensional tasks
In this section, we introduce new parameterizations of the bias force and the indicator function for high-dimensional tasks. Our parameterization is designed around alleviating the problem of sparse training signal, where the model is unable to learn anything due to the failure of collecting meaningful paths that end near the meta-stable state. This problem is especially severe in high-dimensional problems such as large molecules.
Bias force parameterization. To frequently sample the meaningful paths, we parameterize the bias force to guarantee (locally) decreasing the distance between the current molecular state and the target meta-stable state. This is done by parameterizing the bias force as positive atom-wise scaling of the force minimizing the distance, which implies that the bias force can be decomposed into two terms: (1) the direction to decrease the Euclidean distance and (2) the direction that leaves the Euclidean distance invariant. The bias force also satisfies roto-translational equivariance to the current state $\bm{X}_{t}$, aligning with the symmetry of the transition path sampling problem for better generalization.
To be specific, we use a roto-translation $\rho_{t}\in SE(3)$ to align $\bm{R}_{\mathcal{B}}$ with $\bm{R}_{t}$, as we do for indicator function in Equation 2. To achieve $SE(3)$ equivariant, we parameterize the bias force as follows:

### v5

@@ Abstract
Understanding transition pathways between two meta-stable states of a molecular system is crucial to advance drug discovery and material design. However, unbiased molecular dynamics (MD) simulations are computationally infeasible because of the high energy barriers that separate these states. Although recent machine learning techniques are proposed to sample rare events, they are often limited to simple systems and rely on collective variables (CVs) derived from costly domain expertise. In this paper, we introduce a novel approach that trains diffusion path samplers (DPS) to address the transition path sampling (TPS) problem without requiring CVs. We reformulate the problem as an amortized sampling from the transition path distribution by minimizing the log-variance divergence between the path distribution induced by DPS and the transition path distribution. Based on the log-variance divergence, we propose learnable control variates to reduce the variance of gradient estimators and the off-policy training objective with replay buffers and simulated annealing techniques to improve sample efficiency and diversity. We also propose a scale-based equivariant parameterization of the bias forces to ensure scalability for large systems. We extensively evaluate our approach, termed TPS-DPS, on a synthetic system, small peptide, and challenging fast-folding proteins, demonstrating that it produces more realistic and diverse transition pathways than existing baselines. We provide links to our project page and code.
footnotetext: Work done in POSTECH
@@ 1 Introduction
In drug discovery and material design, it is crucial to understand the mechanisms of transitions between meta-stable states of molecular systems, such as protein folding (Salsbury Jr, 2010; Piana et al., 2012), chemical reaction (Mulholland, 2005; Ahn et al., 2019), and nucleation (Thanh et al., 2014; Beaupere et al., 2018). Molecular dynamics (MD) simulations have become one of the most widely used tools for sampling these transitions. However, sampling transition paths with unbiased MD simulations is computationally expensive due to high energy barriers, which cause an exponential decay in the probability of making a transition (Pechukas, 1981).
@@
To address this problem, researchers have developed enhanced sampling approaches such as steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999), umbrella sampling (Torrie & Valleau, 1977; Kästner, 2011), meta-dynamics (Ensing et al., 2006; Branduardi et al., 2012; Bussi & Branduardi, 2015), adaptive biasing force (ABF; Comer et al., 2015), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020) methods. These methods utilize bias forces to facilitate transitions across high energy barriers. They require collective variables (CVs), functions of atomic coordinates designed to capture the slow degree of freedom. Although effective for simple systems, the reliance on expensive domain knowledge limits the applicability of the methods to complex systems where CVs are less understood.
@@
Recently, machine learning has emerged as a promising paradigm for CV-free transition path sampling (TPS) (Das et al., 2021; Lelièvre et al., 2023; Holdijk et al., 2024). The key idea is to parameterize the bias force using a neural network and train it to sample transition paths directly via the biased MD simulation. In particular, Lelièvre et al. (2023) considered reinforcement learning to sample paths escaping meta-stable states. Das et al. (2021); Hua et al. (2024); Holdijk et al. (2024) considered TPS problem as minimizing the reverse Kullback-Leibler (KL) divergence between the path measures induced by the biased MD and the target path measure. However, minimizing the reverse KL divergence suffers from mode collapse, capturing only a subset of modes of the target distribution (Vargas et al., 2023; Richter & Berner, 2024). Furthermore, Das et al. (2021); Lelièvre et al. (2023); Hua et al. (2024); Holdijk et al. (2024) limited their evaluation to low-dimensional synthetic systems or small peptides. Developing machine learning algorithms that generate accurate and diverse transition pathways for complex molecular systems remains an open challenge.
Contribution. In this work, we propose the diffusion path sampler (DPS) to solve the transition path sampling problem. [^1] We coin our method diffusion path sampler since it samples paths using diffusion SDE, similar to diffusion samplers (Zhang & Chen, 2022; Vargas et al., 2023) that use diffusion SDEs for sampling the final state. Our approach, coined TPS-DPS, (1) trains the bias force by minimizing a recently proposed log-variance divergence (Nüsken & Richter, 2021) between the path measure induced by the biased MD and the target path measure, and (2) uses scale-based parameterization of the bias force to handle large systems including fast-folding proteins. Specifically, to leverage desirable properties of the log-variance divergence, such as robustness of gradient estimator and degree of freedom in reference path measure, we propose to learn a control variate for reducing the variance of gradient estimators and employ off-policy training scheme with replay buffer and simulated annealing to improve sample efficiency and prevent the mode collapse.
@@
We also introduce a new scale-based equivariant parameterization for the bias force to frequently sample meaningful paths in training. Our key idea is to predict the atom-wise positive scaling factor of displacement from current molecular states to the target meta-stable state. This guarantees the bias force to decrease the distance between them for every MD step. We also use the Kabsch algorithm (Kabsch, 1976) to align the current molecular states with the target meta-stable state, guaranteeing $SE(3)$ equivariance of bias force for better generalization across the states.
@@
We extensively evaluate our method on the synthetic double-well potential with dual channels, Alanine Dipeptide, and four fast-folding proteins: Chignolin, Trp-cage, BBA, and BBL (Lindorff-Larsen et al., 2011). We compare TPS-DPS with the ML approach (PIPS; Holdijk et al., 2024), as well as classical non-ML methods, e.g., steered MD (SMD; Schlitter et al., 1994; Izrailev et al., 1999). Our experiments demonstrate that TPS-DPS consistently generates realistic and diverse transition paths, similar to the ground truth ensemble. In addition, we do ablation studies of the proposed components.
@@ 2 Related work
Transition path sampling without ML. Metadynamics (Branduardi et al., 2012), on-the-fly probability-enhanced sampling (OPES; Invernizzi & Parrinello, 2020), adaptive biasing force (ABF; Comer et al., 2015), and steered molecular dynamics (SMD; Schlitter et al., 1994; Izrailev et al., 1999) were introduced to explore molecular conformations that are difficult to access by unbiased molecular dynamics (MD) within limited simulation times (Hénin et al., 2022). However, they mostly rely on collective variables (CVs) for high-dimensional problems and are inapplicable to systems with unknown CVs. To sample transition paths without CVs, Dellago et al. (1998) proposed shooting methods that use the Markov chain Monte Carlo (MCMC) procedure on path space. However, these methods often suffer from long mixing times in path space, which hinders the exploration of diverse transition paths. Additionally, high rejection rates can further reduce their efficiency.
@@
Data driven ML approaches. Recently, generative models have been used to sample new transition paths given a dataset of transition paths. Petersen et al. (2023); Triplett & Lu (2023) and Lelièvre et al. (2023) applied diffusion probabilistic models (Ho et al., 2020) and variational auto-encoders (Kingma & Welling, 2013) for transition path sampling, respectively. However, these methods are limited to small systems and struggle to collect data using unbiased MD simulations due to the high energy barriers. Klein et al. (2024); Schreiner et al. (2024); Jing et al. (2024) proposed to accelerate MD using time-coarsened dynamics, but the time-coarsened dynamics cannot capture the details of the transition, e.g., the transition states. Duan et al. (2023); Kim et al. (2024) used neural networks to generate transition states of a given chemical reaction, but cannot generate transition paths.
@@
Data free ML approaches. Without a previously collected dataset, Das et al. (2021); Lelièvre et al. (2023); Sipka et al. (2023); Hua et al. (2024); Holdijk et al. (2024) trained the bias forces to sample transition paths using the biased MD. Lelièvre et al. (2023) used reinforcement learning to train the bias forces but focused on escaping an initial meta-stable state rather than targeting a given meta-stable state. Sipka et al. (2023) used differentiable biased MD simulation to train bias potential and introduce partial back-propagation and graph mini-batching techniques to resolve computational issues in differentiable simulation. Das et al. (2021); Hua et al. (2024); Holdijk et al. (2024) considered the TPS problem as minimizing the reverse KL divergence between path distribution from biased MD and transition path distribution. Das et al. (2021); Hua et al. (2024) limited their evaluation to low-dimensional synthetic systems. In this work, we mainly compare our method with (PIPS; Holdijk et al., 2024). Concurrent to our work, Du et al. (2024) considered the TPS problem as minimizing Doob’s Lagrangian objective with boundary constraints. They parameterized marginal distribution as (mixture) Gaussian path distribution to satisfy the boundary constraints without simulation in training time and sampled transition paths with the bias force derived from the Fokker-Planck equation in inference time.
@@ 3 Transition path sampling with diffusion path samplers
In this section, we introduce our method, coined transition path sampling with diffusion path sampler (TPS-DPS). Our main idea is to formulate the transition path sampling (TPS) problem as a minimization of the log-variance divergence (Nüsken & Richter, 2021) between two path measures: the path measure induced by DPS and that of transition paths. Our main methodological contribution is twofold: (1) a new off-policy training algorithm that minimizes the log-variance divergence with the learnable control variate, replay buffer, and simulated annealing (2) a $SE(3)$ equivariant scale-based parameterization of the bias force that has an inductive bias for dense training signals in large systems.
Figure 1: Problem setup. The sampled transition path (yellow dotted lines) from the state $\mathcal{A}$ to the state $\mathcal{B}$ on the free energy landscape of Alanine Dipeptide. We visualize the snapshots (white circles) of the transition path and the transition state (white star).
@@ 3.1 Problem setup
Our goal is to sample transition paths from one meta-stable state to another meta-stable state given a molecular system. We provide an example of the problem for Alanine Dipeptide in Figure 1. We view this as a task to sample paths from an unbiased molecular dynamics (MD) in Equation 1 conditioned on its starting and ending points of initial and target meta-stable states, respectively. To solve this task, we train the bias force parameterized by a neural network to amortize the sampling procedure.
Molecular dynamics. We consider a MD simulation on time interval $[0,T]$, i.e., the motion of a molecular state $\bm{X}_{t}=(\bm{R}_{t},\bm{V}_{t})\in\mathbb{R}^{6N}$ at time $t$ where $N$ is the number of atoms, $\bm{R}_{t}\in\mathbb{R}^{3N}$ is the atom-wise positions and $\bm{V}_{t}\in\mathbb{R}^{3N}$ is the atom-wise velocities. In particular, we adopt Langevin dynamics (Bussi & Parrinello, 2007) defined as the following SDE:
@@
where $U$, $\bm{m}$, $\gamma$, $k_{B}$, $\lambda$, and $\bm{W}_{t}$ denote the potential energy function, the atom-wise masses, the friction term, the Boltzmann constant, the absolute temperature, and the Brownian motion, respectively, and $\bm{\zeta}\in\mathbb{R}^{3N}$ is a vector of positive infinitesimal values. MD in Equation 1 induces the path measure, denoted by $\mathbb{P}_{0}$, which refers to the positive measure defined on measurable subsets of the path space $\mathcal{C}([0,T];\mathbb{R}^{6N})$ consisting of continuous functions $\bm{X}:[0,T]\rightarrow\mathbb{R}^{6N}$. The path (probability) measure $\mathbb{P}_{0}$ induced by MD assigns high probability to a set of the probable paths when solving MD.
@@
To be specific, we aim to sample from the target path measure $\mathbb{Q}$, which is obtained by reweighting the path measure $\mathbb{P}_{0}$ with the (normalized) indicator function. The indicator function assigns zero weight to paths that do not reach the vicinity of the target position $\bm{R}_{\mathcal{B}}$. Formally, the reweighting function is called the Radon–Nikodym derivative defined as follows:
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}_{0}}(\bm{X})=\frac{1_{\mathcal{B}}(\bm{X})}{Z},\quad 1_{\mathcal{B}}(\bm{X})=\begin{cases}1&\text{if}~{}\lVert\rho_{T}\cdot\bm{R}_{\mathcal{B}}-\bm{R}_{T}\rVert\leq\delta,\\ 0&\text{otherwise},\end{cases}\quad Z=\mathbb{E}_{\mathbb{P}_{0}}\left[1_{\mathcal{B}}(\bm{X})\right],$ (2)
where $\cdot$ denotes group action associated with the $SE(3)$ space and $\rho_{T}\cdot\bm{R}_{\mathcal{B}}$ is the aligned target position by the optimal roto-translation $\rho_{T}\in SE(3)$ to minimize its Euclidean distance to $\bm{R}_{T}$, i.e., $\rho_{T}=\operatorname{argmin}_{\rho\in SE(3)}\lVert\bm{R}_{T}-\rho\cdot\bm{R}_{\mathcal{B}}\rVert$. Such a transformation can be obtained from the Kabsch algorithm in $O(N)$ complexity (Kabsch, 1976).
Note that one may consider naïve rejection sampling to sample transition paths, based on running unbiased MD to sample a path $\bm{X}$ from the path measure $\mathbb{P}_{0}$ and accepting if the path $\bm{X}$ arrives at the neighborhood of the position $\bm{R}_{\mathcal{B}}$ with the radius $\delta$. However, this method struggles to sample transition paths of systems with high energy barriers since the sampled path by unbiased MD rarely reaches the target states, i.e., the rejection ratio is too high.
@@ 3.2 Log-variance minimization
In this section, we propose our algorithm to amortize transition path sampling. Our key idea is to train a neural network to induce a path measure that matches the target path measure $\mathbb{Q}$, using the log-variance divergence between the path measures. We propose a new training scheme to minimize the log-variance divergence based on learning the control variate of its gradient and a replay buffer to improve sample efficiency and diversity.
@@
The first three terms in Equation 6 correspond to the deviation of the biased MD from the unbiased MD integrated over the path sampled from $\mathbb{P}_{\tilde{\bm{v}}}$. The last term reweights the unbiased MD to the target path measure $\mathbb{Q}$. As a result, minimizing Equation 5 could be thought as minimizing the variation between $\mathbb{P}_{\bm{v}_{\theta}}$ and $\mathbb{Q}$. We provide the full derivation in Section A.1. Compared to KL divergence, the log-variance divergence provides a robust gradient estimator and avoids differentiating through the SDE solver (Richter et al., 2020; Nüsken & Richter, 2021).
Minimizing with learnable control variate. To minimize the log-variance divergence, we consider the following loss that replaces the estimation of $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ by learning a scalar parameter $w$:
@@
where $w$ is a control variate that controls the variance of the gradient estimator of $\nabla_{\theta}\mathcal{L}(\theta,w)$ without changing the gradient. Note that we set $\tilde{\bm{v}}=\bm{v}_{\theta}$ in Equation 6, which implies that the gradient of Equation 7 coincides with the KL divergence (Richter et al., 2020; Nüsken & Richter, 2021). When optimized, the control variate $w$ estimates the expectation $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ since $\operatornamewithlimits{argmin}_{w}\mathcal{L}(\theta,w)=\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$. Thus, jointly optimizing $(\theta,w)$ with the gradient step can be interpreted as jointly minimizing log-variance divergence and estimating $\mathbb{E}_{\mathbb{P}_{\bm{v}_{\theta}}}[F_{\bm{v}_{\theta},\bm{v}_{\theta}}]$ utilizing $w$.
Off-policy training with replay buffer and simulated annealing. To leverage the degree of freedom in reference path measure for the log-variance divergence, we allow discrepancy between reference path measure and current path measure, called off-policy training, which is widely used in discrete-time reinforcement learning (Mnih et al., 2013; Bengio et al., 2021). For the sample efficiency, we reuse the samples with a replay buffer $\mathcal{D}$ which stores path samples from the path measure $\mathbb{P}_{\bm{v}_{\bar{\theta}}}$ associated with previous policies $\bm{v}_{\bar{\theta}}$. Our modified loss function $\mathcal{L}^{\mathcal{D}}$ with $\mathcal{D}$ is defined as follows:
@@
The replay buffer also prevents mode collapse, using diverse paths from different path measures. Additionally, in line with other off-policy training algorithms (Malkin et al., 2022; Kim et al., 2023), we utilize the simulated annealing technique to sample diverse paths that cross high-energy barriers.
@@
Algorithm 1 Training
2:  for $i=1,\ldots,I$ do
5:     for $j=1,\ldots,J$ do
6:        Sample $K$ data $\{\bm{x}_{0:L}^{(k)}\}^{K}_{k=1}$ from $\hat{\mathcal{D}}$.
7:        Update $\theta$ and $w$ with the gradient of $\frac{1}{K}\sum_{k=1}^{K}\left(\log\frac{p_{0}(\bm{x}^{(k)}_{0:L})1_{\mathcal{B}}(\bm{x}^{(k)}_{0:L})}{p_{\bm{v}_{\theta}}(\bm{x}^{(k)}_{0:L})}-w\right)^{2}$.
8:     end for
9:  end for
We describe our training algorithm in Algorithm 1. Overall, our off-policy training algorithm iterates through four steps: (1) sampling paths from the biased MD simulation with current policy $\bm{v}_{\theta}$ at high temperature, (2) storing sampled paths in the replay buffer $\hat{\mathcal{D}}$, (3) sampling a batch of the paths from the replay buffer, and (4) training current policy $\bm{v}_{\theta}$ by minimizing the loss in Equation 10. After training, the biased MD simulation can directly sample transition paths from the target path measure.
3.3 Parameterization for large systems
In this section, we introduce new parameterizations of the bias force and the indicator function to sample transition paths of large systems. Our parameterization is designed to alleviate the problem of sparse training signal, where the model struggles to collect meaningful paths that end at the vicinity of the target meta-stable state in training. This problem is especially severe in large systems.
Bias force parameterization. To frequently sample the meaningful paths, we aim to parameterize the bias force which guarantees to reduce the distance between the current molecular state and the target meta-stable state for every MD step. This is achieved by predicting the atom-wise positive scaling factor of the direction to the aligned target meta-stable state rather than predicting force or potential directly. Moreover, we design the bias force to satisfy roto-translational equivariance to the current molecular state input $\bm{X}_{t}$, aligning with the symmetry of the transition path sampling problem for better generalization and parameter efficiency.
To be specific, we achieve the inductive bias for dense training signals and the $SE(3)$ equivariance to the current molecular state input $\bm{X}_{t}$ by parameterizing the bias force as follows: